OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -3) and a directrix at y = 3.

OpenStudy (mertsj):

The parabola has vertex (0,0) and is concave downward.

OpenStudy (mertsj):

p=3

OpenStudy (mertsj):

x^2=4py

OpenStudy (anonymous):

I got y2 = -3x

OpenStudy (anonymous):

Is it correct?

OpenStudy (mertsj):

x^2=4py and p = 3 Replace p with 3

OpenStudy (anonymous):

y2 = -12x, I am sure of it.

OpenStudy (anonymous):

Thank You :)

OpenStudy (mertsj):

The equation you have given opens to the left. The parabola you described opens down. Therefore the x term is the one that is squared. Draw the graph of y^2=-12x http://www.wolframalpha.com/input/?i=graph+y^2%3D-12x

OpenStudy (mertsj):

See that is not the equation of a parabola whose directrix is a horizontal line.